# Optimized View and Geometry Distillation from Multi-view Diffuser

### 17 Dec 2023 v2

@@ 4.2 Evaluation Protocol
| SyncDreamer [19] | 0.0261 | 0.5421 |
| Wonder3D [21] | 0.0199 | 0.6244 |
| Ours | 0.0240 | 0.5688 |
@@ 4.3 Ablation Study
We validate our design choices by ablating 4 major model variants, that are without Stable Diffusion(SD) unconditional predicated noise, without input view supervise, without the first and second stages of DreamBooth(DB). The texture color of the 3D models generated by without SD unconditional noise exhibit bias, and the shapes are not smooth. Not using input view supervision will result in the inability to recover the same texture details as the input image, and not using DreamBooth will result in blurring of texture details.
@@ 4.5 Single View Reconstruction
We evaluate the quality of the reconstructed geometry of different methods. The quantitative results are summarized in Table 1, and the qualitative comparisons are presented in Fig. 7. The quality of Wonder3D [21] shape reconstruction depends on the perspective of the input view, such as Train shown in Fig. 7, where Wonder3D [21] generated incorrect prediction results. The shape generated by SyncDreamer [19] undergoes deformation due to the camera pose in the input view. One-2-3-45 [16] attempts to reconstruct meshes from the multiview-inconsistent outputs of Zero-1-to-3 [17]. While it can capture coarse geometries, it loses important details in the process. In contrast, our method can achieve good reconstruction quality and texture in terms of geometric structure and texture. In the paper, all the surface extraction demonstrated by our method is built on the instant-NGP based SDF reconstruction method [9], and we use Blender Cycles [3] to render the results.
@@ 4.6 Novel View Synthesis
| Method | PSNR | SSIM | LPIPS |
| Realfusion [22] | 15.26 | 0.722 | 0.283 |
| Zero-1-to-3 [17] | 18.93 | 0.779 | 0.166 |
| SyncDreamer [19] | 20.05 | 0.798 | 0.146 |
| Wonder3D [21] | 26.07 | 0.924 | 0.065 |
| Ours | 23.62 | 0.931 | 0.075 |
We evaluate the quality of novel view synthesis for different methods. The quantitative results are presented in Table 2, and the qualitative results can be found in Fig. 4. Zero-1-to-3 [17] produces visually reasonable images, but they lack multi-view consistency since it operates on each view independently (we don’t show the results of Zero-1-to-3). Although SyncDreamer [19] introduces a volume attention scheme to enhance the consistency of multi-view images, their model is sensitive to the elevation degrees of the input images and tends to produce unreasonable results. Wonder3D ensures 3D consistency by generating normal maps, but may result in incorrect results for some input viewpoints.
@@ 5 Conclusions
In this work, we introduce an optimized approach for distilling geometry and views from a multi-view diffuser, with a specific focus on the Zero-1-to-3 model. We observed that the direct application of the SDS/SJC technique to Zero-1-to-3 is often suboptimal, primarily due to bias issues inherent in unconditional noise. To address this, we propose an Unbiased Score Distillation (USD) strategy by leveraging unconditioned noises from a 2D diffusion model to effectively enhance the optimized radiance field. Moreover, we developed a two-stage DreamBooth refinement process to improve the rendering of views. This process ensures consistency across multiple perspectives while simultaneously enhancing image quality.
While we have identified and addressed the bias issue in the Zero-1-to-3 model, the underlying causes remain to be fully understood. Future research will delve into the theoretical aspects of this bias problem. Additionally, we aim to explore the potential applications of USD in other fields, such as image translation and view synthesis.

### 08 Mar 2024 v3

@@ 4.2 Evaluation Protocol
| SyncDreamer [23] (NeuS) | 0.0261 | 0.5421 |
| Wonder3D [24] (iNGP+NeuS) | 0.0199 | 0.6244 |
| Ours (NeuS) | 0.0240 | 0.5688 |
| Ours (iNGP+NeuS) | 0.0177 | 0.6330 |
@@ 4.3 Ablation Study
We validate our design choices by ablating 4 major model variants, that are without Unbiased Score Distillation, without reference view distillation loss, without the two-stage DreamBooth(DB). The texture color of the 3D models geneDirated by without USD exhibit bias, and the shapes are not smooth. Not using reference view supervision will result in the inability to recover the same texture details as the input image, and not using DB will result in blurring of texture details.
@@ 4.5 Single View Reconstruction
We evaluate the quality of the reconstructed geometry of different methods. The quantitative results are summarized in Table 3, and the qualitative comparisons are presented in Fig. 9. The quality of Wonder3D [24] shape reconstruction depends on the perspective of the input view, such as ‘Train’ shown in Fig. 9, where Wonder3D generated incorrect prediction results. The shape generated by SyncDreamer [23] undergoes deformation due to the camera pose in the input view. One-2-3-45 [20] attempts to reconstruct meshes from the multiview-inconsistent outputs of Zero-1-to-3 [21]. While it can capture coarse geometries, it loses important details in the process. In contrast, our method can achieve good reconstruction quality and texture in terms of geometric structure and texture. In the paper, all the surface extraction demonstrated by our method is built on the Instant NGP(iNGP) [28] based SDF reconstruction method [10], and we use Blender Cycles [3] to render the results.
@@ 4.6 Novel View Synthesis
We evaluate the quality of novel view synthesis for different methods. The quantitative results are presented in Table 3, and the qualitative results can be found in Fig. 4. Zero-1-to-3 [21] produces visually reasonable images, but they lack multi-view consistency since it operates on each view independently (we don’t show the results of Zero-1-to-3). Although SyncDreamer [23] introduces a volume attention scheme to enhance the consistency of multi-view images, their model is sensitive to the elevation degrees of the input images and tends to produce unreasonable results. Wonder3D [24] ensures 3D consistency by generating normal maps, but may result in incorrect results for some input viewpoints.
@@ 4.7 Text-to-image-to-3D
As a case study, we combine text-to-image models, i.e., the Stable Diffusion or Imagen [36] to generate 3D models from text. We show some examples in Fig. 12. Compared to DreamFusion [30] and ProlificDreamer [48], our method shows no multi-face Janus problem and conforms to the text faithfully.
@@ 5 Conclusions
In this work, we introduce an optimized approach for distilling geometry and views from a multi-view diffuser, with a specific focus on the Zero-1-to-3 model. We observed that the direct application of the SDS/SJC technique to Zero-1-to-3 is often suboptimal, primarily due to bias issues inherent in unconditional noise. To address this, we propose an Unbiased Score Distillation (USD) strategy by leveraging unconditioned noises from a 2D diffusion model to effectively enhance the optimized radiance field. Moreover, we developed a two-stage DreamBooth refinement process to improve the rendering of views. This process ensures consistency across multiple perspectives while simultaneously enhancing image quality. While we have identified and addressed the bias issue in the Zero-1-to-3 model, the underlying causes remain to be fully understood. Future research will delve into the theoretical aspects of this bias problem. Additionally, we aim to explore the potential applications of USD in other fields, such as image translation and view synthesis.